OpenStudy (jeanfifa09):

A bus company charges $15 per person for the first 50 passengers who buy tickets for a gold outing. The company offers a discount of $0.10 per ticket for each passenger in excess of 50. What number of passengers will provide maximum income for the bus company ?

jabez177 (jabez177):

Ehhhh... It's been a while since I've seen these questions.

jabez177 (jabez177):

@pooja195 @Luigi0210 @hartnn

OpenStudy (jeanfifa09):

I need help doing it , I'm a little bit stuck

OpenStudy (shaik0124):

what have u done so far

OpenStudy (shaik0124):

for first 50 passengers ticket cost will be $15 then from 51st passenget onwards the ticket cost will be $14.9

OpenStudy (jeanfifa09):

I need to find the number of passengers that will provide maximum income for the bus company

OpenStudy (jeanfifa09):

This is an optimization problem ,for calc 1

ganeshie8 (ganeshie8):

is the complete statement of the problem ?

OpenStudy (jeanfifa09):

I have I = # of passengers * price per passengers I = (50+x)(10-.10x) I = 500-5x+10x-0.1x^2 Do I take the derivative ? I'm stuck right there ,

OpenStudy (jeanfifa09):

That is the complete statement .

OpenStudy (jeanfifa09):

What do you mean by complete statement ?

OpenStudy (jeanfifa09):

I meant to write above I = (50+x)(15-0.1x)

Miracrown (miracrown):

yes - we need derivatives now, what is the derivative of this function?

OpenStudy (jeanfifa09):

dI/dx=10-0.2X

Miracrown (miracrown):

Sorry, my connection was acting fuzzy. So yes if you want to optimize you need to derive and make this equal to 0

OpenStudy (jeanfifa09):

0=10-0.2x ? 0.2x=10 X= 50 ?

Miracrown (miracrown):

Let me show you: dI/dx= 5-(0.1)(2)x

Miracrown (miracrown):

I = 500-5x+10x-0.1x^2 =5x-(0.1)(2)x

Miracrown (miracrown):

you see?

Miracrown (miracrown):

So this new one, could you equal it to 0 and find the x?

OpenStudy (jeanfifa09):

I made a mistake originally , I fixed it to be : I= (50+x)(15-0.10x) Which equals to I = 750+10x-0.1x^2 I wrote (50+x)(10-0.10x) which was wrong . The price is $15 not $10

Miracrown (miracrown):

Oh, ok. So the derivative would be?

OpenStudy (jeanfifa09):

The derrative of that would be dI/dx= 10-0.2X

Miracrown (miracrown):

Yep! So when its equal to zero?

OpenStudy (jeanfifa09):

0=10-.2x .2x=10 X=50?

Miracrown (miracrown):

Yess!!!

Miracrown (miracrown):

so at 50 persons the price get to its maximum :)

OpenStudy (jeanfifa09):

After what do I do next ? Plug it in the original equation ?

Miracrown (miracrown):

Nope! That's your answer X = 50 !

OpenStudy (jeanfifa09):

Oh ok

OpenStudy (jeanfifa09):

Thank you so much!!!

Miracrown (miracrown):

You're most welcome! Good luck with your studies! :D